OpenStudy (anonymous):

find two integers whose sum is 17 and whose product is 60

jhonyy9 (jhonyy9):

a+b=17 a*b=60

jhonyy9 (jhonyy9):

a+b=17 --- subtract from both sides b a+b-b=17-b a=17-b so this value of a you can subtirute in place of a inside second equation than what will get ?

OpenStudy (ladiesman217):

its 12 & 5

jhonyy9 (jhonyy9):

ok but how you have got these ? step by step please

OpenStudy (anonymous):

thank you @jhonyy9

OpenStudy (ladiesman217):

i thought about a clock for some reason

OpenStudy (anonymous):

Let one no=x then 2nd no=17-x x(17-x)=60 17x-x^2=60 x^2-17x+60=0 x^2-12x-5x+60=0 x(x-12)-5(x-12)=0 (x-12)(x-5)=0 x-12=0 gives x=12 &2nd no =17-12=5 x-5=0 gives x=5 & 2nd no=17-5=12 nos are 5& 12